Initial Setup at Front Bent Gregorian - 17 September 2011 UT

Observer(s): J. Hill (Tucson)
Telescope Operator: S. Allanson (LBTO)
Telescope Support: K. Newton (LBTO)
Instrument Support: T. Shih (LBTO)
Software Support: N. Cushing (LBTO)


We opened late because of humidity. We confirmed the RFBG results from last night. We were closed by clouds and humidity before we found any star on LFBG.

Humidty kept us closed for the rest of the night. No new progress was made on LFBG. Andrew and John share the opinion that the strong coma seen last night must be from the swing arm being out of position. Therefore we want to reverse the change made last week. (See the email to telescopework)


There was a full power outage during the day (MGIO power system maintenance), so EVERYTHING has been reinitialized. Thanks to Kevin, Steve, Norm, et al. for getting everything restarted.

DXSMIRTCFCollimation.dat was updated with the following global offsets:
X=-0.611 Y=-1.662 Z=+0.738 RX=78.3 RY=336.4

Details (times in UT)

Humidity is high, so opening is delayed. Steve is fighting with communication between GCS and AzCam (on both sides). azcamserver apparently requires multiple reboots to cure its problems for AGW2. GCSL was waiting for PCS to acknowlege the coordinated acquisition offset in binocular mode while GCSR failed to find a closed-dome star.


04:10 Begin manually unlatching doors

04:20 Open with cirrus clouds in sky.


04:27 Preset M5_ ACQUIRE - found the ghost next to the bright star on the initial pointing (star within ~20 arcsec). Image has some mild astigmatism and defocus. Center with IE=19 CA=15.

04:35 Preset ACTIVE - Seeing on guider is about 1 arcsec FWHM. EL=59

04:40 Preset ACTIVE - About 4 arcsec of astigmatism at EL=38, pointing within 5 arcsec of the previous star.

The updates to the DX M2 collimation model are confirmed.

IRTC on both sides

04:49 Synchronized ACQUIRE presets to the M5_ star. Dead on the star on the right, no star on left.

Update IE/CA on left to match the right side. SX M2 Xglobal to +7.5 mm.

We failed to get the IRTC pointing model on the left side. but applying it doesn't find the star.


05:04 Changing the authorization from IRTC on both sides to IRTC on left and NONE on right loaded the LUCI SX pointing model on the right side. And it did not reload the pointing model terms on the left. ?????

05:08 ACQUIRE preset on left side. - ptspiral (moving the mount in monocular mode)

Why did telescope just start to slew???? Steve clicked azhold to stop it at 05:11:28. See Issue 3641.

05:15 New preset to same target (M5_1568) causes unwrap from 38 to 356. ??? Something is goofy here........

05:20 Kill PCS and restart.


05:22 ACQUIRE preset on right side.

The fresh PCS still wants to unwrap to AZ=348. IE=18 CA=19. Star is here, so we were pointing at some unknown place at 05:10.


This time both pointing models went to LUCIFERSX - this may be the proper behaviour.
Changed them manually to IRTC pointing models.

Heavier clouds are rolling naked eye stars visible. So we close.


05:36 Close for clouds.

-- JohnHill - 17 Sep 2011
Topic revision: r3 - 17 Sep 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback